OpenStudy (anonymous):

Given the quadratic function f(x) = x2 – 12x + 36, find a value of x such that f(x) = 9.

OpenStudy (anonymous):

They're telling you that f(x) = x^2 - 12x + 36, and they're also tellig you that f(x) = 9 so we can rewrite this to become x^2 - 12x + 36 = 9 this becomes x^2 - 12x +27 = 0 and then you use your normal solving methods to find x.

OpenStudy (anonymous):

so wouldnt the outcome be f(9)=9?

OpenStudy (gw2011):

x^2-12x+36=9 x^2-12x+27=0 (x-9)(x-3)=0 x-9=0 x=9 x-3=0 x=3 Therefore, the solution is x=9 and x=3 You prove this by substitution: x^2-12x+36 (9)^2-12(9)+36 81-108+36=9 (3)^2-12(3)+36 9-36+36=9